# Towards Evaluating Generative Audio: Insights from Neural Audio Codec Embedding Distances

### v1

@@ 1 Introduction
This paper presents a systematic study of audio quality prediction using FAD and MMD across three encoder domains: EnC, DAC, and a novel enhanced DAC (DACe) trained with synthetic tonal data. While FAD and MMD are typically used for zero-shot, reference-free evaluation, their efficacy fundamentally depends on whether the underlying embeddings capture human perceptual judgments. To validate this foundational requirement, we first evaluate these metrics on the simpler task of predicting full-bandwidth audio quality given a reference signal and comparing correlations with MUSHRA [13] scores. This controlled, signal-to-signal comparison serves as a necessary validation step: if an embedding space does not reflect human perception in this setting, it is unlikely to succeed in a true reference-free scenario. Furthermore, since high-quality subjective datasets for generative audio are scarce (e.g., MusicEval [14], DCASE 2023 Task 7 [15], and SongEval [16] are either not full bandwidth or contain MP3 artifacts), and the choice of reference set is crucial [3], our approach of direct signal-to-signal comparisons against a reliable MUSHRA test reduces uncertainty and enables clearer conclusions.
@@ 2 Distance Measures
$\text{FAD}^{2}(\mathbf{X},\mathbf{Y})=\|\mu_{\mathbf{X}}-\mu_{\mathbf{Y}}\|_{2}^{2}+\text{tr}\left(\Sigma_{\mathbf{X}}+\Sigma_{\mathbf{Y}}-2\sqrt{\Sigma_{\mathbf{X}}\Sigma_{\mathbf{Y}}}\right).$
This definition assumes a multivariate Gaussian distribution, is sensitive to dataset sample size, and incurs high computational costs. Although $\mathrm{FAD}_{\infty}$ [3] mitigates sample size bias by extrapolating to infinite samples, other limitations remain. To address these issues, a distribution-free, unbiased, and computationally efficient metric based on MMD [2] is employed.
@@
$\text{MMD}^{2}(P,Q)$ $=\mathbb{E}_{x,x^{\prime}}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}}[k(y,y^{\prime})]$ (2)
@@
$\widehat{\mathrm{MMD}}^{2}=$ $\frac{1}{n(n-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{m(m-1)}\sum_{i\neq j}k(y_{i},y_{j})$ (3)
@@
$\text{MMD}_{s}=\alpha\cdot\widehat{\text{MMD}}^{2},$
@@ 3 Neural Audio Codecs (NACs)
Descript Audio Codec (DAC): The DAC targets general audio coding at 44.1 kHz. We evaluated the publicly available 16 kb/s model [20] using a MUSHRA listening test (Figure 1), which confirmed that DAC outperforms EnC. A stereo listening test was conducted with 15 stereo excerpts (48 kHz) from the ODAQ set [21]. Music excerpts were retained, while six dialogue-heavy excerpts were excluded to ensure a fair comparison with EnC. DAC was tested at 16 and 32 kb/s, encoding the left and right channels independently (i.e., dual-mono coding). Input signals were resampled to 44.1 kHz for coding and back to 48 kHz for listening tests. Results show that DAC at 16 kb/s outperformed EnC at 24 kb/s. However, DAC struggles with tonal content, such as glockenspiel, a limitation also reported by its authors [7].
@@
Enhanced Descript Audio Codec (DACe): We developed DACe by improving DAC and training it with up to 32 10-bit codebooks (30 kb/s, 48 kHz model). DACe was trained using the same hyperparameters as DAC, but with a batch size of 48 instead of 72, on a diverse 720-hour, 48 kHz dataset [22] covering 18 music genres, speech, and isolated instruments. To address DAC’s weakness on tonal material, we added 32 hours of real tonal audio and, more critically, generated synthetic tonal signals on the fly during training to expose the time-domain codec to challenging content. Each synthetic tonal excerpt was created by dynamically simulating multiple tonal events per sample, with frequency, amplitude, and decay evolving according to randomized onset timing and Poisson-distributed event density. Balanced sampling ensured that 33% of each mini-batch consisted exclusively of challenging synthetic tonal excerpts. A MUSHRA test (Figure 2) compared DAC and DACe at 32 kb/s stereo using dual-mono coding. Test excerpts were drawn from an internal dataset (since ODAQ excerpts were not sufficiently challenging), with none overlapping the training set, and were selected based on maximal Mel-loss [7] deviation between the two NACs. As shown in Figure 2, DACe consistently outperforms DAC, with strong gains on tonal items (e.g., triangles and the MPEG glockenspiel test file sm02) and on a panned speech excerpt. Overall, DACe outperforms DAC, and both surpass EnC.
@@ 4.1 FAD and MMD in NAC embedding domain
Between the two distance measures, we observe that FAD correlates more strongly with subjective perceptual scores. From Eqs. (1)–(4), the squared Fréchet Audio Distance, $\text{FAD}^{2}(\mathbf{X},\mathbf{Y})$, depends solely on the empirical means and covariances of the embeddings. In contrast, the unbiased MMD estimator, $\widehat{\text{MMD}}^{2}$, is a U-statistic sensitive to higher-order moments via a kernel $k$ [17].

### v2

@@ 1 Introduction
This paper presents a systematic study of audio quality prediction using FAD and MMD across three encoder domains: EnC, DAC, and a novel enhanced DAC (DACe) trained with synthetic tonal data. While FAD and MMD are typically used for zero-shot, reference-free evaluation, their efficacy fundamentally depends on whether the underlying embeddings capture human perceptual judgments. To validate this foundational requirement, we first evaluate these metrics on the simpler task of predicting full-bandwidth audio quality given a reference signal and comparing correlations with MUSHRA [22] scores. This controlled, signal-to-signal comparison serves as a necessary validation step: if an embedding space does not reflect human perception in this setting, it is unlikely to succeed in a true reference-free scenario. Furthermore, since high-quality subjective datasets for generative audio are scarce (e.g., MusicEval [19], DCASE 2023 Task 7 [2], and SongEval [31] are either not full-bandwidth or contain MP3 artifacts), and the choice of reference set is crucial [14], our approach of direct signal-to-signal comparisons against a reliable MUSHRA test reduces uncertainty and enables clearer conclusions.
@@ 2 Distance Measures
$\mathbf{FAD}^{2}(\mathbf{X},\mathbf{Y})=\|\mu_{\mathbf{X}}-\mu_{\mathbf{Y}}\|_{2}^{2}+\text{tr}\left(\Sigma_{\mathbf{X}}+\Sigma_{\mathbf{Y}}-2\sqrt{\Sigma_{\mathbf{X}}\Sigma_{\mathbf{Y}}}\right).$
This definition assumes a multivariate Gaussian distribution, is sensitive to dataset sample size, and incurs high computational costs. Although $\mathbf{FAD}_{\infty}$ [14] mitigates sample size bias by extrapolating to infinite samples, other limitations remain. To address these issues, a distribution-free, unbiased, and computationally efficient metric based on MMD [16] is employed.
@@
$\mathbf{MMD}^{2}(P,Q)$ $=\mathbb{E}_{x,x^{\prime}}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}}[k(y,y^{\prime})]$ (2)
@@
$\widehat{\mathbf{MMD}}^{2}=$ $\frac{1}{n(n-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{m(m-1)}\sum_{i\neq j}k(y_{i},y_{j})$ (3)
@@
$\mathbf{MMD}_{s}=\alpha\cdot\widehat{\mathbf{MMD}}^{2},$
@@ 3 Neural Audio Codecs (NACs)
Descript Audio Codec (DAC): The DAC targets general audio coding at 44.1 kHz. We evaluated the publicly available 16 kb/s model [7] using a MUSHRA listening test (Fig. 2), which confirmed that DAC outperforms EnC. A stereo listening test was conducted over headphones with 15 stereo excerpts (48 kHz) from the ODAQ set [28]. Music excerpts were retained, while six dialogue-heavy excerpts were excluded to ensure a fair comparison with EnC. DAC was tested at 16 and 32 kb/s, encoding the left and right channels independently (i.e., dual-mono coding). Input signals were resampled to 44.1 kHz for coding and back to 48 kHz for listening tests. Results show that DAC at 16 kb/s outperformed EnC at 24 kb/s. However, DAC struggles with tonal content, such as glockenspiel, a limitation also reported by its authors [18].
@@
Enhanced Descript Audio Codec (DACe): We developed DACe by improving DAC and training it with up to 32 10-bit codebooks, corresponding to $\sim$30 kb/s per channel at 48 kHz. DACe was trained using the same hyperparameters as DAC, but with a batch size of 48 instead of 72, on a diverse 720-hour, 48 kHz mono dataset [5] covering 18 music genres, speech, and isolated instruments. To address DAC’s weakness on tonal material, we added 32 hours of real tonal audio and, more critically, generated synthetic tonal signals on the fly during training to expose the time-domain codec to challenging content. Each synthetic tonal excerpt was created by dynamically simulating multiple tonal events per sample, with frequency, amplitude, and decay evolving according to randomized onset timing and Poisson-distributed event density. Balanced sampling ensured that 33% of each mini-batch consisted exclusively of challenging synthetic tonal excerpts. A MUSHRA test (Fig. 2) compared DAC and DACe at 32 kb/s stereo using dual-mono coding. Test excerpts were drawn from an internal dataset (since ODAQ excerpts were not sufficiently challenging), with none overlapping the training set, and were selected based on maximal Mel-loss [18] deviation between the two NACs. As shown in Fig. 2, DACe consistently outperforms DAC, with strong gains on tonal items (e.g., triangles and the MPEG glockenspiel test file sm02) and on a panned speech excerpt. Overall, DACe outperforms DAC, and both surpass EnC.
@@ 4.1 FAD and MMD in NAC embedding domain
Between the two distance measures, we observe that FAD correlates more strongly with subjective perceptual scores. From Eqs. (1)–(4), the squared Fréchet Audio Distance, $\mathbf{FAD}^{2}(\mathbf{X},\mathbf{Y})$, depends solely on the empirical means and covariances of the embeddings. In contrast, the unbiased MMD estimator, $\widehat{\mathbf{MMD}}^{2}$, is a U-statistic sensitive to higher-order moments via a kernel $k$ [12].